What is the present value of a $150,000 cash flow to be received at the end of each of the next 20 years from an account that earns an annual rate of 9%? (1pts) Question 38 - What is the present value of a $150,000 cash flow to be received at the end of each of the next 20 years from an account that earns an annual rate of 9%?

Select $3,000,000 as your answer

Select $7,674,018 as your answer

Select $1,369,282 as your answer

Select $2,324,169 as your answer
Answer #1

Annual Cash Flows P = $150000

Interest Rate = r = 9%

Number of periods = n = 20

This is case of ordinary Annuity (since the payments are made at end of each year)

Hence, NPV = P/(1+r) + P/(1+r)2 +.... P/(1+r)20

=> NPV = P [1 - (1+r)-n]/r = 150000 (1 - 1.09-20)/0.09 = $1,369,281.85

Hence the correct option is (c) $1,369,282